### v2

@@ 6.5 Critical Decision Points with Suboptimal Data
Figure 8: Illustration of the impact of varying the number of parameters and dataset usage on state-action novelty estimation. The error bar represents the standard deviation of state-action novelty across 25 trajectories.
@@ 7 Conclusion
We present NBDI, an approach for learning terminated skills through a state-action novelty module that leverages offline, task-agnostic datasets. Our approach significantly outperforms previous baselines in solving complex, long-horizon tasks and shows effectiveness even under significant changes in environment configuration of downstream tasks. A promising direction for future work is to use novelty-based decision point identification to learn variable-length skills in meta-reinforcement learning (Nam et al., 2022).
@@ Appendix E Train/transfer Domain Similarity
To apply our method to downstream environments with significantly different overall layouts (e.g., a maze with a completely new structure or a larger-scale block stacking environment with more blocks in random positions), it is necessary to collect cropped image centered around the agent from task-agnostic demonstrations (visualizations of cropped images are available in Figure 2). This way, we can extract consistent structural information that can be extracted across multiple task-agnostic trajectories, which then can be used to detect state-action novelty based decision points even in downstream tasks with significantly different environment configurations. On the other hand, for environments where the overall layout remains consistent (e.g. the positions of manipulatable objects in the kitchen environment do not change during downstream task), we can use any state information that sufficiently describes the environment to apply our method (see Appendix F for environment details).
@@
Appendix F Data and Environment Details
We evaluate NBDI on three environments: one simulated navigation task (Maze navigation) and two simulated robotic manipulation tasks (Kitchen and Sparse block stacking). We employ the environment configuration and dataset provided by Pertsch et al. (2021a) (Maze navigation and Sparse block stacking) and Fu et al. (2021) (Kitchen) (CC BY 4.0). Note that task and environment setup differ between the training data and the downstream task, demonstrating the model’s capacity to handle unseen downstream tasks.
@@ Appendix F Data and Environment Details
Differences to Pertsch et al. (2021a).
While Pertsch et al. (2021a) employed a block stacking environment with dense rewards (the agent is rewarded based on the height of the stacked tower and for actions like picking up or lifting blocks), we evaluated our model and the baselines in a sparse block stacking environment, leading to different performance outcomes. In this setting, the agent is rewarded solely for the height of the tower it constructs, which increases the task complexity. Furthermore, there have been consistent reports indicating that performance in large maze environments displays a high level of sensitivity to random seeds, primarily due to the high stochasticity of the task. For the five different seeds that we used to compare the algorithms, we found the performance to be generally lower than what was previously reported, mainly due to the sensitivity to seeds.
Appendix G Implementation Details
Our codebase builds upon the released code of SPiRL (Pertsch et al., 2021a). Our code is available at: https://anonymous.4open.science/r/nbdi-4D44
G.1 Termination Improvement and Novelty
We present termination improvement and conditional action novelty in a simple $8\times 8$ grid maze domain. As we mentioned in Appendix F, training data for the state-action novelty module is collected with diverse tasks. Thus, we set up the environment as follows: We randomly select one starting location and three goal locations to generate three trajectories for each goal location. The goal-reaching data collection policy randomly executes a discrete action, moving towards four different directions (left, right, forward, backward) while avoiding moving toward walls. The agent receives a binary reward when reaching the goal state.
@@ G.1 Termination Improvement and Novelty
G.2 State-Action Novelty Module
@@ G.3 BC+SAC and IQL+SAC
G.4 LOVE and Relative novelty
@@
G.5 Experiments Compute Resources

### v3

@@ 6.6 Model Capacity and Critical Decision Point Detection
7 Applying NBDI to Skill-based Meta Reinforcement Learning
To explore the broader applicability of our approach, we investigated its compatibility with a skill-based approach, SiMPL (Nam et al., 2022) that leverages task-agnostic demonstrations to solve challenging long-horizon, sparse-reward meta-RL tasks. Specifically, we apply our method during the skill extraction phase to learn variable-length skills in place of fixed-length skills. In the maze environment, we randomly sampled 10 goal locations for meta-training. The table below compares the success rate of meta-policies trained with SiMPL (fixed-length skills) and our method during the meta-training phase across episodes. In Figure 9, the result shows that the extracted variable-length skills allows the meta-policy to better promote knowledge transfer between different tasks, helping the meta-policy in combining the skills to complete complex tasks. We report mean success rates on the 10 meta-training tasks across 3 different seeds with standard errors. Furthermore, during the target task learning phase, the meta-policy learned through our approach leads to significantly better sample efficiency on the unseen target task. Results in Table 2 indicate that our approach can be effectively integrated with a broader class of skill-based methods that leverage task-agnostic demonstrations.
ep20 ep100 ep300 ep500
SiMPL 0.143$\pm$0.063 0.593$\pm$0.191 0.667$\pm$0.193 0.990$\pm$0.006
SiMPL+NBDI 0.560$\pm$0.121 0.960$\pm$0.021 0.980$\pm$0.011 0.993$\pm$0.003
Table 2: Comparison of success rates in the maze environment during the target task learning phase of SiMPL with standard errors over three random seeds.
@@ 8 Conclusion
We present NBDI, an approach for learning terminated skills through a state-action novelty module that leverages offline, task-agnostic datasets. Our approach significantly outperforms previous baselines in solving complex, long-horizon tasks and shows effectiveness even under significant changes in environment configuration of downstream tasks. A promising direction for future work is to use novelty-based decision point identification to learn variable-length skills in offline skill execution (Ajay et al., 2020; Hakhamaneshi et al., 2021).
Acknowledgments
This work was partly supported by Institute of Information &\&& communications Technology Planning &\&& Evaluation (IITP) grant funded by the Korea government (MSIT) (No. RS-2022-II220311, Development of Goal-Oriented Reinforcement Learning Techniques for Contact-Rich Robotic Manipulation of Everyday Objects, No. RS-2024-00457882, AI Research Hub Project, and No. RS-2019-II190079, Artificial Intelligence Graduate School Program (Korea University)), the IITP(Institute of Information &\&& Coummunications Technology Planning &\&& Evaluation)-ITRC(Information Technology Research Center) grant funded by the Korea government(Ministry of Science and ICT)(IITP-2025-RS-2024-00436857), the NRF (RS-2024-00451162) funded by the Ministry of Science and ICT, Korea, BK21 Four project of the National Research Foundation of Korea, and the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT)(RS-2025-00560367), and the IITP under the Artificial Intelligence Star Fellowship support program to nurture the best talents (IITP-2025-RS-2025-02304828) grant funded by the Korea government (MSIT).
Impact Statement
This paper presents work whose goal is to advance the field of Reinforcement Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ B.2 Reinforcement Learning with NBDI
B.3 Threshold for environments
In practice, the thresholds we tuned through experiments (see Appendix A.1) approximately correspond to the 97th percentile of the novelty values computed over task-agnostic demonstrations—e.g., kitchen = 0.3 (97.47th percentile), maze = 50 (96.86th percentile), and block stacking = 40 (96.12th percentile). We hope the percentile-based guidance would help the readers more easily apply our method across new environments.
@@ Appendix E Train/transfer Domain Similarity
To apply our method to downstream environments with significantly different overall layouts (e.g., a maze with a completely new structure or a larger-scale block stacking environment with more blocks in random positions), it is necessary to collect cropped image centered around the agent from task-agnostic demonstrations (visualizations of cropped images are available in Figure 2). This way, we can extract consistent structural information that can be extracted across multiple task-agnostic trajectories, which then can be used to detect state-action novelty based decision points even in downstream tasks with significantly different environment configurations. On the other hand, for environments where the overall layout remains consistent (e.g. the positions of manipulatable objects in the kitchen environment do not change during downstream task), we can use any state information that sufficiently describes the environment to apply our method (see Appendix H for environment details).
@@
Appendix F Task-Agnostic Dataset Size and Coverage
The number of task-agnostic trajectories collected for each environment were as follows: maze (85,000 trajectories), block stacking (37,000 trajectories), and kitchen (400 trajectories). Since we used image-based observations in the maze and block stacking environments, and the downstream tasks involve significant configuration changes (e.g., entirely new maze layouts or a larger-scale block stacking environment with more blocks in random positions), we require a larger set of task-agnostic demonstrations compared to the kitchen environment, where the overall layout remains consistent across tasks.
Since our goal in maze and block stacking environment is to solve downstream tasks with significantly different environment configuration, we do not assume that task-agnostic demonstrations fully cover the state space of the downstream tasks. However, it is important that they provide a good coverage of the observation space. Note we are using cropped image centered around the agent from task-agnostic demonstrations as observations. This way, we can extract consistent structural information that can be extracted across multiple task-agnostic trajectories, which then can be used to detect state-action novelty based decision points even in downstream tasks with significantly different environment configurations.
Appendix G Comparison to baseline methods on discrete state/action space
To provide an intuitive comparison between NBDI, Relative Novelty (Şimşek & Barto, 2004), and LOVE (Jiang et al., 2022) in discrete settings, we conducted a case study using a grid-based maze environment (Figure 18). This allowed us to directly visualize and compare termination points across methods in a controlled, discrete domain. To align with our task-agnostic setup, we collected diverse expert-level demonstrations from random start and goal positions, and used these datasets to extract skills and termination points.
Relative Novelty defines novelty at a given state as the ratio between the average novelty of future and past states, measured within a fixed-size sliding window (n_lag). A high value indicates that the agent transitions from a familiar region to a less familiar one. Following the original formulation, we only evaluated states with sufficient trajectory context on both sides of the window. As shown in the visualization, Relative Novelty is highly dependent on transition history, often identifying only a subset of bottleneck states.
LOVE uses a variational inference framework to extract skills based on the Minimum Description Length (MDL) principle in that its objective is to ”effectively compress a sequence of data by factoring out common structure”. While LOVE employs a variational inference framework to implement the MDL principle, we used Byte Pair Encoding (BPE) to extract skills in discrete setting as BPE is a specific formulation of MDL (Gallé, 2019) and it offers a more intuitive and interpretable formulation of compression. Termination points were visualized based on where the segmented skills terminated during 100 goal-reaching tasks. The results show that terminations vary significantly with the number of trajectories used to extract skills, as LOVE focuses on capturing common structure rather than consistent bottlenecks.
NBDI terminates skills based on both conditional action novelty and state novelty (Section 4.1). The visualization demonstrates that NBDI consistently identifies key bottleneck states and exhibits robustness to the number of trajectories collected in contrast to other methods. Moreover, states with high state novelty often correspond to regions that are rare or hard to reach within the task-agnostic dataset. By increasing the decision frequency in such unfamiliar states, NBDI promotes more effective exploration of the state space.
Figure 18: Visualization of skill termination points in a grid-based maze environment using NBDI (Conditional action novelty, State novelty), Relative Novelty, and LOVE (Minimum Description Length). Darker shades represent states with a relatively higher frequency of termination occurrences. Black cells are impassable wall regions in the environment. Top: termination points computed from 100 trajectories. Bottom: results from 10,000 trajectories.
Appendix H Data and Environment Details
We evaluate NBDI on three environments: one simulated navigation task (Maze navigation) and two simulated robotic manipulation tasks (Kitchen and Sparse block stacking). We employ the environment configuration and dataset provided by (Pertsch et al., 2021a) (Maze navigation and Sparse block stacking) and (Fu et al., 2021) (Kitchen) (CC BY 4.0). Note that task and environment setup differ between the training data and the downstream task, demonstrating the model’s capacity to handle unseen downstream tasks.
@@ Appendix H Data and Environment Details
Differences to (Pertsch et al., 2021a).
While (Pertsch et al., 2021a) employed a block stacking environment with dense rewards (the agent is rewarded based on the height of the stacked tower and for actions like picking up or lifting blocks), we evaluated our model and the baselines in a sparse block stacking environment, leading to different performance outcomes. In this setting, the agent is rewarded solely for the height of the tower it constructs, which increases the task complexity. Furthermore, there have been consistent reports indicating that performance in large maze environments displays a high level of sensitivity to random seeds, primarily due to the high stochasticity of the task. For the five different seeds that we used to compare the algorithms, we found the performance to be generally lower than what was previously reported, mainly due to the sensitivity to seeds.
Appendix I Implementation Details
Our codebase builds upon the released code of SPiRL (Pertsch et al., 2021a). Our code is available at: https://github.com/ku-dmlab/NBDI.
I.1 Termination Improvement and Novelty
We present termination improvement and conditional action novelty in a simple $8\times 8$ grid maze domain. As we mentioned in Appendix H, training data for the state-action novelty module is collected with diverse tasks. Thus, we set up the environment as follows: We randomly select one starting location and three goal locations to generate three trajectories for each goal location. The goal-reaching data collection policy randomly executes a discrete action, moving towards four different directions (left, right, forward, backward) while avoiding moving toward walls. The agent receives a binary reward when reaching the goal state.
@@ I.1 Termination Improvement and Novelty
I.2 State-Action Novelty Module
@@ I.3 BC+SAC and IQL+SAC
I.4 LOVE and Relative novelty
@@
I.5 Experiments Compute Resources